3t-7/15=13/15 I need to show my work
Answer by stanbon(75887)

3t-7/15=13/15
-------------------
Add 7/15 to both sides:
3t = 20/15
--------
Reduce the fraction:
3t = 4/3
---
Divide both sides by 3 to get:
t = 4/9
=============
Cheers,
Stan H.
=============